# BRcal: An R Package to Boldness-Recalibrate Probability Predictions

### v2

@@ 3 The BRcal package
3.1 Overview
The main contributions of the BRcal package are (i) boldness-recalibration, (ii) MLE-recalibration, (iii) calibration assessment, and (iv) supporting visualizations. The brcal() function facilitates boldness-recalibration at a user-specified level of calibration, The mle_recal() function facilitates MLE-recalibration. The bayes_ms() and llo_lrt() functions facilitate Bayesian and frequentist approaches to calibration assessment, respectively. The plot_params() function provides a visualization for the constrained optimization via a color contour plot of the posterior model probability in (12). The lineplot() function provides a visualization for examining how individual predictions change under boldness-recalibration and MLE-recalibration (optionally). The package also provides a supporting function, LLO(), for LLO-adjusting via any specified parameters.
The BRcal package has three main dependencies on other R packages. The brcal function requires the nloptr package for optimization [27]. Additionally, the plotting functions make use of the fields [28] and ggplot2 packages [29].
@@ 3.2 Data
The BRcal package includes two built-in datasets called foreclosure and hockey. The foreclosure dataset includes probability predictions of foreclosure and the foreclosure status of 5,000 randomly selected housing transactions in 2010 from Wayne County, Michigan. These data arose from a project that developed a monitoring strategy for the housing market using a spatially-adjusted model of foreclosure risk [30]. The probability of foreclosure was estimated using a mixture model of spatial kernel density estimates. The model was trained during the first six months of 2005 using data from Wayne County, Michigan, which is a pre-Great Recession time period. The foreclosure dataframe includes three columns:
  • y - integer; sale type, 1 = foreclosure, 0 = regular sale
  • x - numeric; predicted probabilities of foreclosure
  • year - numeric; year of observed foreclosure or regular sale
The hockey dataset includes probability predictions of a home team win from FiveThirtyEight [31] and the winner of each game in the 2020-2021 National Hockey League (NHL) season. These data were used in [1], and more information about the data can be found there. The hockey data frame includes four columns:
@@ 3.3 The brcal() Function
The core function of the BRcal package is the brcal() function, which implements boldness-recalibration. It takes two required arguments, x a vector of probability predictions, and y a vector of binary event outcomes. The other key argument is t, which represents the minimum desired probability of calibration as seen in (9) and is set to 0.95 by default. Users should adjust this setting based on the level of risk they are willing to accept by sacrificing some calibration in favor of increased boldness. The arguments for brcal() are summarized in Table 1.
@@
The brcal() function uses the optimization capabilities of nloptr() for the nonlinear constrained optimization to find $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$ and optim() to find $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$. We primarily focus on how brcal() uses the nloptr() function in this section. Discussion of optim() to find the MLEs, which are used in boldness-recalibration, can be found in Section 3.4. The nloptr package implements both global and local optimization routines and both gradient-based and derivative-free approaches. For all sets of predicted probabilities the authors encountered or generated, the constraint surface (the posterior model probability) has been unimodal. Within the constrained region, the objective surface (the standard deviation) has also been unimodal. With this in mind, we chose to focus on local gradient based approaches of which there were three options that supported non-linear inequality constraints: Sequential Least-Squares Quadratic Programming (SLSQP) [32, 33], Augmented Lagrangian algorithm (AUGLAG) [34, 35], and Method of Moving Asymptotes (MMA) [36]. Of these, AUGLAG produced the most stable results in terms of convergence in boldness-recalibration problems and is set as the default optimization routine in brcal().
The implementation of AUGLAG provided by nloptr reformulates the constrained optimization problem as a two-stage optimization. The inner stage optimizes the objective function with a penalty for violating the constraint rather than use the constraint directly. The outer stage governs the penalty. If the inner stage returns a value that violates the constraint despite the penalty, the outer stage imposes a larger penalty. Note that since the inner stage is itself an unconstrained optimization routine, users may specify any optimization routine that supports nonlinear objective functions. For the purpose of this package, we found that SLSQP as the inner optimization routine in AUGLAG provided the most stable results, thus SLSQP is set as the default inner optimization routine.
@@
As with any optimization problem, it is important to verify convergence and adjust arguments as needed to achieve convergence. The brcal() function provides both of these by allowing, but not requiring, full adjustment of the nloptr() and optim() routines, and returning important convergence information. The default arguments used in brcal() have been effective in most applications explored by the authors. A few options for stopping criteria, starting values, parameter bounds, and level of output are specified directly in brcal(). All other nloptr() arguments can be optionally adjusted by passing them as a list to the opts argument. Similarly, all arguments in optim() can be adjusted by passing them as a list to optim_options.
@@
To improve convergence, we set the default starting location of the nloptr() optimizer to the MLEs for the original probability predictions. The constraint function in (12) is maximized at the MLEs, and since the constraint surface has been empirically unimodal, these MLEs lie at the center of the constraint region. This is illustrated in Figure 1 by the $\times$ denoting the MLEs located at the center of the shaded region. By starting at the center of the plausible region of parameter values, this prevents the optimizer from being unable to locate the region of interest. We allow users to toggle this option via the start_at_MLEs parameter. By default, this parameter is set to TRUE. If set to FALSE, users can specify an alternative starting value via argument x0. Should users specify their own x0, the authors recommend imposing bounds on the parameter space using lb and ub to set lower and upper bounds. respectively, on $\delta$ and $\gamma$.
@@ 3.4 Calibration Assessment
While the goal of boldness-recalibration is to embolden when predictions are accurate, users can first assess calibration using the bayes_ms() and llo_lrt() functions. Both functions provide MLEs for $\delta$ and $\gamma$. The bayes_ms() function uses the Bayesian model selection-based approach described in Section 2.1. In addition to MLEs, this function returns the posterior model probability of calibration calculated via (3), and the BIC for both $M_{c}$ and $M_{u}$ in (6) and (7). The llo_lrt() function conducts the likelihood ratio test-based approach to assess calibration, details can be found in [1]. In addition to MLEs, this function returns the p-value associated with the test, and the test statistic. Details of the arguments for bayes_ms(), llo_lrt(), and other supporting functions in the BRcal package can be found in Table 2.
@@
Both bayes_ms() and llo_lrt() rely on a call to the optim() function to maximize the log of the likelihood in (1) to find $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$. By default optim() minimizes the function it is given. To follow this convention, we choose to minimize the negative of the log likelihood. Additionally, optim() offers several options for optimization routines. Given that parameter space of $\gamma$ is unbounded but $\delta>0$, we choose to re-parameterize the log likelihood for sake of optimization. Instead of operating on $\delta$, we use $\tau=log(\delta)\in\mathbb{R}$ [23]. With this version of the log likelihood, we choose to use the default routine, "Nelder-Mead" [37]. While by default we operate on $\tau$ instead of $\delta$ in the optimization step, bayes_ms() and llo_lrt() also have a parameter called tau which will change the output of the function to be in terms of $\tau$ instead of $\delta$. To make any adjustments to the call to optim(), users can pass any arguments to optim() using ....
3.5 MLE-Recalibration
@@
Should users want to LLO-adjust based on parameter values other than the MLEs or $(100*t)\%$ boldness-recalibration parameters, we also provide the LLO() function. This function takes argument x as the vector of probabilities to which the LLO function will be applied. Arguments delta and gamma correspond to $\delta$ and $\gamma$, respectively, in (2). The returned vector of probabilities is $\mathbf{x}^{\prime}$ where $x^{\prime}_{i}=c(x_{i};\delta^{\prime},\gamma^{\prime}$), and $\mathbf{x},\delta^{\prime},\gamma^{\prime}$ are the user provided arguments.
@@ 3.6 Visualization Functions
The BRcal package provides two types of visualizations to conduct and assess the impact of boldness-recalibration: (i) a lineplot visualizing how individual predictions change under various LLO-adjustments and (ii) a color image and contour plot of the posterior model probability as a function of $\delta$ and $\gamma$. Examples of both are provided in Figures 2 and 3 in Section 4. This section describes the implementation details for these plot types. Description of the plotting function arguments and default values can be found in Table 3.
@@ 3.6.1 Lineplot
The first visualization provided is a lineplot showing how probability predictions x change after LLO-adjustment implemented in the lineplot() function and utilizes ggplot2 graphics [29]. This function allows users to pass a set of predictions x and outcomes y, and specify which predictions (original, MLE-recalibrated, tboldness-recalibrated) they are interested in plotting. By default, only the original and MLE-recalibrated probabilities are plotted. To omit either the original or MLE-recalibrated sets, users can set plot_original or plot_MLE to FALSE. To specify levels of boldness-recalibration, users can pass a vector of values for $t$ via t_levels. The resulting plot shows each specified set of probabilities side-by-side with lines connecting corresponding predictions. See Figure 2 for an example. The lines are color coded based on the corresponding outcome, where blue represents $y_{i}$ = 1 (i.e. an “event”) and red represents $y_{i}$ = 0 (i.e. a “non-event”). The goal is to visualize (i) the overall change in boldness between the sets (e.g. spread increases/decreases under $95\%$ boldness-recalibration), (ii) how individual predictions change (e.g. a prediction of 0.7 is adjusted to 0.8), and (iii) if predictions with high predicted probabilities typically result in events (e.g. point/lines for higher probabilities are typically blue, indicating an “event”).
@@
When the number of predictions for your lineplot is very large, plotting may be slow. To ease the graphical burden, we offer three thinning options to reduce the amount of predictions that are visualized. The thin_to argument allows users to specify some $n_{thin}$ = the number of predictions they would like to reduce their sample to. In this case, lineplot() will randomly select $n_{thin}$ predictions from x and their corresponding outcomes in y. Using thin_prop allows users to select the proportion of predictions to plot, which results in a random selection of that proportion size. Lastly, specifying the thin_by argument to some number $m$ where lineplot() then selects every $m$ observation from the predictions. Note that under all thinning strategies, all underlying calculations including the posterior model probabilities, boldness-recalibration parameters, and MLEs are conducted using the full set of probability predictions. Thinning only reduces the number of observations seen on the plot.
@@
Once users have settled on the desired LLO-adjustments to be plotted (i.e. desired levels of boldness-recalibration, MLE-recalibration, etc), you can specify return_df=TRUE to return the underlying dataframe of adjusted probabilities to avoid recalculation of these values. This dataframe can then be reused by passing it to argument df, indicating to lineplot() that no recalculations are needed. For example, it is useful to reuse df so that users do not need to wait for the underlying boldness-recalibration calculations each time they make minor cosmetic adjustments to the plot, like modifying axes labels or font sizes.

### v3

@@ 4 The BRcal package
The main contributions of the BRcal package are (i) boldness-recalibration, (ii) MLE recalibration, (iii) calibration assessment, and (iv) supporting visualizations. The brcal() function conducts boldness-recalibration at a user-specified level of calibration. The mle_recal() function facilitates MLE recalibration. The bayes_ms() and llo_lrt() functions performs Bayesian and frequentist approaches to calibration assessment, respectively. The plot_params() function provides a visualization for the constrained optimization via a color contour plot of the posterior model probability in (13). The lineplot() function provides a visualization for examining how individual predictions change under boldness-recalibration and MLE recalibration (optionally). The package also provides a supporting function, LLO(), for LLO-adjusting via any specified parameters.
The BRcal package has three main dependencies on other R packages. The brcal() function requires the nloptr package, which is the R implementation of the NLopt library, for nonlinear constrained optimization [29]. Additionally, the plotting functions make use of the fields [35] and ggplot2 packages [36].
@@ 4.1 Data
The BRcal package includes two built-in datasets called hockey and foreclosure.
The hockey dataset includes probability predictions of a home team win from FiveThirtyEight [37] and the winner of each game in the 2020-2021 National Hockey League (NHL) season. These data were used in [1], and more information about the data can be found there. All package demonstrations in this section utilize this data. The hockey data frame includes four columns:
@@
The foreclosure dataset includes probability predictions of foreclosure and the foreclosure status of 5,000 randomly selected housing transactions in 2010 from Wayne County, Michigan. These data arose from a project that developed a monitoring strategy for the housing market using a spatially-adjusted model of foreclosure risk [38]. The probability of foreclosure was estimated using a mixture model of spatial kernel density estimates. The model was trained during the first six months of 2005 using data from Wayne County, Michigan, which is a pre-Great Recession time period. A case study analysis of this data is presented in Section 5. The foreclosure dataframe includes three columns:
  • y - integer; sale type, 1 = foreclosure, 0 = regular sale
  • x - numeric; predicted probabilities of foreclosure
  • year - numeric; year of observed foreclosure or regular sale
@@ 4.2 The brcal() Function
The core function of the BRcal package is the brcal() function, which implements boldness-recalibration using the nonlinear constrained optimization described in the previous section. By default, brcal() performs 95% boldness-recalibration (i.e., $t$=0.95). Users can select a different minimum desired probability of calibration as seen in (9) using argument t. Users should adjust this setting based on the level of risk they are willing to accept by sacrificing some calibration in favor of increased boldness. The arguments for brcal() are summarized in Table 1.
@@
Based on the discussion in Section 3.2.1, the nloptr() function uses the local gradient-based implementation of Auglag, with SLSQP as the inner routine, by default to find $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$ in brcal(). To leverage the closed-form derivative information, we supply the Jacobians of our objective and constraint functions from Appendices A.3 and A.4 to nloptr() within brcal(). The brcal() function also uses optim() to find $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$, however we primarily focus on how brcal() uses the nloptr() function in this section. Discussion of optim() is reserved for Section 4.3.
By default, brcal() prints each step of the optimization process from nloptr(). To suppress all output from nloptr(), users can set print_level=0. Other options to minimize output can be found in the nloptr() documentation and can be passed to brcal() via print_level. Below we show an example where the full output from the first and last steps are shown, but the in-between stages are truncated by ... for readability.
> br95 <- brcal(hockey\$x, hockey\$y)
iteration: 1
Ψx = (0.945397, 1.400573)
Ψf(x) = -0.123926
Ψg(x) = -0.048849
iteration: 146
Ψx = (0.872885, 1.958574)
Ψf(x) = -0.165339
Ψg(x) = -0.000000
> br95
nloptr
Call:
nloptr::nloptr(x0 = c(0.945397, 1.400573), eval_f = obj_f, eval_grad_f = obj_grad_f,
    eval_g_ineq = constr_g, eval_jac_g_ineq = constr_grad_g,
    opts = list(algorithm = "NLOPT_LD_AUGLAG", maxeval = 500,
           maxtime = -1, xtol_rel = 1e-06, print_level = 3,
           local_opts = list(algorithm = "NLOPT_LD_SLSQP",
             eval_grad_f = obj_grad_f, eval_jac_g_ineq = constr_grad_g,
             xtol_rel = 1e-06)), t = 0.95, tau = FALSE, Pmc = 0.5,
    epsilon = 2.22044604925031e-16)
Minimization using NLopt version 2.7.1
NLopt solver status: 4 ( NLOPT_XTOL_REACHED: Optimization stopped because
xtol_rel or xtol_abs (above) was reached. )
Number of Iterations....: 146
Termination conditions:  maxeval: 500Ψxtol_rel: 1e-06
Number of inequality constraints:  1
Number of equality constraints:    0
Optimal value of objective function:  -0.165338762728956
Optimal value of controls: 0.8728849 1.958574
\$Pmc
[1] 0.5
\$t
[1] 0.95
\$BR_params
[1] 0.8728849 1.9585745
\$sb
[1] 0.1653388
\$probs
   [1] 0.5741797 0.7322471 0.5734560 0.5020176 0.5799966 0.3786775
   [7] 0.3228442 0.4505939 0.5180415 0.2241145 0.5229929 0.5265926
 [865] 0.8054830 0.7286748 0.2542999 0.4597869
Notice the large amount of output from nloptr() which is returned by brcal() in nloptr. As with any optimization problem, it is important to verify convergence and adjust arguments as needed to achieve convergence. We include this copious output to allow users to check that the routine converged properly. While this output shows the achieved boldness-recalibration estimates and standard deviation, we choose to return these separately in the list items BR_params and sb, respectively, for easier extraction. The resulting set of boldness-recalibrated probabilities are also returned for easier extraction.
@@
The brcal() function also provides, but does not require, full adjustment of the nloptr() and optim() routines. The default arguments used in brcal() have been effective in most applications explored by the authors. A few options for stopping criteria, starting values, parameter bounds, and level of output are specified directly in brcal(). All other nloptr() arguments, including the selection of optimization routine, can be optionally adjusted by passing them as a list to the opts argument. Similarly, all arguments in optim() can be adjusted by passing them as a list to optim_options.
@@
For reasons discussed in Section 3.2.1, we set the default starting location of the nloptr() optimizer to the MLEs for the original probability predictions. We allow users to toggle this option via the start_at_MLEs parameter. By default, this parameter is set to TRUE. If set to FALSE, users can specify an alternative starting value via argument x0. Should users specify their own x0, the authors recommend imposing bounds on the parameter space using lb and ub to set lower and upper bounds, respectively, on $\delta$ and $\gamma$.
@@ 4.3 Calibration Assessment
While the goal of boldness-recalibration is to embolden when predictions are accurate, users can first assess calibration using the bayes_ms() and llo_lrt() functions. The bayes_ms() function uses the Bayesian model selection-based approach described in Section 2. The llo_lrt() function conducts the likelihood ratio test-based approach to assess calibration, details can be found in [1]. Details of the arguments for bayes_ms(), llo_lrt(), and other supporting functions in the BRcal package can be found in Table 2.
@@
Both bayes_ms() and llo_lrt() provide numerical solutions for the MLEs for $\delta$ and $\gamma$ using the optim() function. By default these functions use the "Nelder-Mead" routine as discussed in Section 3.1.1. While results are reported in terms of $\delta$, users can specify tau=TRUE to report results in terms of $\tau$ instead. To make any adjustments to the call to optim(), users can pass any arguments to optim() using ....
Notice from the output below that bayes_ms() returns seven quantities in the form of a list. The prior model probability for $M_{c}$ is returned in Pmc to remind users what value was specified (0.5 by default). List items BIC_Mc and BIC_Mu in (6) and (7) are the BICs for the calibrated model and uncalibrated model, respectively. The Bayes factor comparing $M_{u}$ to $M_{c}$ in (5) is returned in BF and posterior_model_prob is the posterior model probability of calibration from (3). The MLEs for the vector x are returned in MLEs. Additionally, the list returned by the call to optim() to get the MLEs is returned in sublist optim_details.
> bt <- bayes_ms(hockey\$x, hockey\$y)
> bt
\$Pmc
[1] 0.5
\$BIC_Mc
[1] 1148.637
\$BIC_Mu
[1] 1157.902
\$BF
[1] 0.009730538
\$posterior_model_prob
[1] 0.9903632
\$MLEs
[1] 0.9453966 1.4005730
\$optim_details
\$optim_details\$par
[1] 0.9453966 1.4005730
\$optim_details\$value
[1] 572.1846
\$optim_details\$counts
function gradient
      63       NA
\$optim_details\$convergence
[1] 0
\$optim_details\$message
NULL
> llo_lrt(hockey\$x, hockey\$y)
\$test_stat
[1] 4.267411
\$pval
[1] 0.1183977
\$MLEs
[1] 0.9453966 1.4005730
\$optim_details
\$optim_details\$par
[1] 0.9453966 1.4005730
\$optim_details\$value
[1] 572.1846
\$optim_details\$counts
function gradient
      63       NA
\$optim_details\$convergence
[1] 0
\$optim_details\$message
NULL
4.4 MLE recalibration
@@
> mle_recal(hockey\$x, hockey\$y)
\$probs
  [1] 0.5633646 0.6814612 0.5628440 0.5117077 0.5675526 0.4223768
  [7] 0.3802169 0.4748424 0.5231574 0.3000678 0.5266953 0.5292678
[865] 0.7421488 0.6786381 0.3255808 0.4814569
\$MLEs
[1] 0.9453966 1.4005730
\$optim_details
\$optim_details\$par
[1] 0.9453966 1.4005730
\$optim_details\$value
[1] 572.1846
\$optim_details\$counts
function gradient
      63       NA
\$optim_details\$convergence
[1] 0
\$optim_details\$message
NULL
By default, mle_recal() returns a list with (i) the vector of maximally calibrated predictions in probs, (ii) the MLEs in MLEs and (iii) the details from optim() in optim_details. Should a user only want the vector of MLE recalibrated probabilities, they can set the option probs_only = TRUE. An example of when users may want this option is for assessing the calibration of the MLE recalibrated set. By specifying probs_only = TRUE users can pass the MLE recalibrated set directly from mle_recal() to bayes_ms() to get the maximum achievable probability of calibration for that set. If the probabilities can only achieve a low probability of calibration, we would not recommend further emboldening the predictions, as they are likely poorly calibrated.
4.5 LLO-Adjustment
Should users want to LLO-adjust based on parameter values other than the MLEs or $(100*t)\%$ boldness-recalibration parameters, we also provide the LLO() function. The LLO() function allows users to directly LLO-adjust a vector of probability predictions x under any specified shift parameter delta and scale parameter gamma. While most users of this package may not use the LLO() function directly, there are some cases where users may want to do so. One example of this could be when you have MLEs for $\delta$ and $\gamma$ based on historical data and want to LLO-adjust unlabeled out-of-sample predictions. A user could use the LLO function to LLO-adjust the out of sample predictions by passing them to function argument x, and passing the MLEs to arguments delta and gamma. Another use case could be when you have boldness-recalibration parameter estimates from brcal() but did not save the boldness-recalibrated probabilities. Users could LLO-adjust their probabilities using the boldness-recalibration parameter estimates using LLO() without needing to re-compute them using brcal(), thus avoiding unnecessary computation time. This is demonstrated below with the 95% boldness-recalibration estimates saved in Section 4.2. Notice in the output below that we get the same boldness-recalibrated probabilities as we did using brcal() in Section 4.2.
> LLO(hockey\$x, delta=br95\$BR_params[1], gamma=br95\$BR_params[2])
  [1] 0.5741797 0.7322471 0.5734560 0.5020176 0.5799966 0.3786775
  [7] 0.3228442 0.4505939 0.5180415 0.2241145 0.5229929 0.5265926
[865] 0.8054830 0.7286748 0.2542999 0.4597869
@@ 4.6 Visualization Functions
The BRcal package provides two types of visualizations to conduct and assess the impact of boldness-recalibration: (i) a lineplot visualizing how individual predictions change under various LLO-adjustments and (ii) a color image and contour plot of the posterior model probability as a function of $\delta$ and $\gamma$. Examples of both are provided in Figures 5 and 4 in Section 5. This section describes the implementation details for these plot types. Description of the plotting function arguments and default values can be found in Table 3.
@@ 4.6.1 Lineplot
The first visualization provided is a lineplot showing how probability predictions x change after LLO-adjustment implemented in the lineplot() function and utilizes ggplot2 graphics [36]. By default, only the original and MLE recalibrated probabilities are plotted. To omit either the original or MLE recalibrated sets, users can set plot_original or plot_MLE to FALSE. To specify levels of boldness-recalibration, users can pass a vector of values for $t$ via t_levels. The resulting plot shows each specified set of probabilities side-by-side with lines connecting corresponding predictions. The code below generates the plot in Figure 2 as an example. The lines are color coded based on the corresponding outcome, where blue represents $y_{i}$ = 1 (i.e., an “event”) and red represents $y_{i}$ = 0 (i.e., a “non-event”). The goal is to visualize (i) the overall change in boldness between the sets (e.g. spread increases/decreases under $95\%$ boldness-recalibration), (ii) how individual predictions change (e.g. a prediction of 0.7 is adjusted to 0.8), and (iii) if predictions with high predicted probabilities typically result in events (e.g. point/lines for higher probabilities are typically blue, indicating an “event”).
> lineplot(hockey\$x, hockey\$y, t_levels=c(0.95, 0.9, 0.8))
Figure 2: Demonstration of the lineplot() function on the hockey data.
@@
When the number of predictions for your lineplot is very large, plotting may be slow. To ease the graphical burden, we offer three thinning options to reduce the amount of predictions that are visualized. The thin_to argument allows users to specify the number of predictions, $n_{thin}$, they would like to reduce their sample to. In this case, lineplot() will randomly select $n_{thin}$ predictions from x and their corresponding outcomes in y. Using thin_prop allows users to select the proportion of predictions to plot, which results in a random selection of that proportion size. Lastly, specifying the thin_by argument to some number $m$ where lineplot() then selects every $m$ observation from the predictions. Note that under all thinning strategies, all underlying calculations including the posterior model probabilities, boldness-recalibration parameters, and MLEs are conducted using the full set of probability predictions. Thinning only reduces the number of observations seen on the plot.
@@
Once users have settled on the desired LLO-adjustments to be plotted (i.e., desired levels of boldness-recalibration, MLE recalibration, etc), you can specify return_df=TRUE to return the underlying dataframe of adjusted probabilities to avoid recalculation of these values. This dataframe can then be reused by passing it to argument df, indicating to lineplot() that no recalculations are needed. For example, it is useful to reuse df so that users do not need to wait for the underlying boldness-recalibration calculations each time they make minor cosmetic adjustments to the plot, like modifying axes labels or font sizes.